# Structure of Kerr black hole spacetimes in Weyl conformal gravity

### v1

@@ Abstract
We analyze the stationary, uncharged, rotating, vacuum solution to Weyl conformal gravity. We elucidate the causal and ergoregion structure of the spacetimes found in the parameter space of the metric for positive mass. These are then compared to the corollary Kerr, Kerr-de Sitter, and Kerr-Anti-de Sitter solutions to general relativity. Additionally, we investigate and derive the extremal limits for both the horizons and ergosurfaces of the spacetimes.
@@ 1 Introduction
where $G_{\mu\nu}$ is the Einstein curvature tensor, and $T_{\mu\nu}$ is the stress-energy tensor. We use geometrized units, where $G=c=\hbar=1$, and a metric $g_{\mu\nu}$ signature of (,+,+,+)(-,+,+,+)( - , + , + , + ).
@@
Observations of flat galactic rotation curves [10, 11] and anomalies in the dynamics of galaxy clusters [12, 13] have required positing the existence of invisible dark matter. Meanwhile, the discovery of the accelerating expansion of the universe [14] has pointed towards the need for a dark energy component described by a cosmological constant $\Lambda$. The dominant cosmological paradigm, $\Lambda$CDM, includes such additional energy contributions in the Einstein field equations (1).
Serious qualms against the $\Lambda$CDM paradigm have arisen from the seemingly ad hoc nature of the necessitated invocation of dark matter, a lack of direct detections of it [15], and the vacuum energy density attributed to dark energy being $\sim 120$ orders of magnitude lower than what is predicted by particle physics theory [16].
The desire to resolve these tensions motivates work into alternative gravity theories such as Modified Newtonian dynamics (MOND) [17], massive gravity [18], Horndeski gravity [19, 20], and others.
@@
Requiring conformal invariance means that we now have the CG field equations
$W_{\mu\nu}=\frac{1}{4\ \alpha_{\text{g}}}\ T_{\mu\nu},$ (2)
where $W_{\mu\nu}$ is the Bach curvature tensor, and $\alpha_{\text{g}}$ is a dimensionless gravitational coupling constant [23, 24]. While $G_{\mu\nu}$ contains up to second-order derivatives of the metric, $W_{\mu\nu}$ has up to fourth-order derivatives, making CG a fourth-order theory.
@@
[…]
@@
$B(r)=1-\frac{2\widetilde{M}}{r}-3\beta\gamma+\gamma r-\kappa r^{2}.$ (5)
@@
The presence of the linear $\gamma r$ and quadratic $-\kappa r^{2}$ terms in the lapse function allows CG to account for flat galactic rotation curves [29, 30, 31] and cosmological expansion [23, 27, 32] without dark matter or dark energy. Fitting $\gamma$ and $\kappa$ to relevant phenomena and mass distributions remains an active field of research [23, 27, 33, 26, 34]. CG successfully replicates results of GR in such classical tests as perihelion precession and time dilation [35, 36] for small values of $\gamma$ and $\kappa$, making it a viable alternative to GR. It is worth noting that CG has had issues with handling galaxy cluster dynamics [37, 38].
@@
[…]
@@
From this, we recover the GR Kerr metric for $\gamma,\kappa\rightarrow 0$. An appropriate conformal transformation and taking $\gamma\rightarrow 0$ reduces CG Kerr to GR Kerr-de Sitter (Kerr-dS) for $\kappa>0$, and GR Kerr-Anti-de Sitter (Kerr-AdS) for $\kappa<0$. Conformal transformations also show that CG Kerr (6) is equivalent to CG Schwarzschild (3) when $a\rightarrow 0$ [26, 42]. As such a transformation has not been explicitly shown in Boyer-Linquist coordinates in existing literature, we demonstrate this in A.
@@
In this work, we explore the parameter space of the CG parameters $\gamma$ and $\kappa$, the mass $\beta$, and the spin parameter $a$. We consider both positive and negative values of $\gamma$ and $\kappa$. Values of $\gamma<0$ are potentially required to permit gravitational lensing [36], in contrast to $\gamma>0$ values used to fit galaxy rotation curves and cosmological phenomena [23, 27, 33]. Work has been done to reconcile this stark disparity [46].
[…]
@@
We begin by introducing the equations governing the presence of ergosurfaces and horizons in section 2. Afterwards, we establish the means to determine the causal structure and nature of horizons and ergosurfaces in section 3. The relevant equations for the extremal limits of CG Kerr spacetimes are derived in section 4. We then move on to establishing the general structure of CG Kerr spacetimes for $\kappa=0$ along with elucidating the role of $\gamma$ in section 5. The effects of varying $\kappa$ and the spin $a$ are handled in sections 6 and 7 respectively. We summarize our conclusions in section 8.
@@ 2.1 Ergoregions and ergosurfaces
The ergosurfaces, where $g_{tt}=0$, that bound ergoregions are then known as static limits. On the equatorial plane $(\theta=\frac{\uppi}{2})$, this condition is given by
$\Delta^{\mathcal{E}}\equiv\left(\rho^{2}g_{tt}\right)\big{|}_{\theta=\frac{\uppi}{2}}=-kr^{4}+r^{2}-2\widetilde{M}r=0.$ (8)
@@
As a quartic with real coefficients, we may have four, two, or no real roots. Similar to GR, the CG Kerr solution possesses a ring singularity at $(r,\theta)=\left(0,\frac{\uppi}{2}\right)$. While real roots of $\Delta^{\mathcal{E}}$ may be found for $r<0$, we confine our work to discussions of features found for $r\geq 0$. For the most part, only three ergosurfaces are found for $r\geq 0$.
@@ 2.2 Horizons
Horizons or static limits, meanwhile, demarcate where $g_{rr}$ switches sign, and thus where $r$ goes from being a spacelike coordinate to a timelike one. These occur at coordinate singularities of $g_{rr}$, where its denominator $\Delta^{\mathrm{H}}$, defined in (7), vanishes as
@@
In order to better orient ourselves, we review the ordering of ergosurfaces $\mathcal{E}$ and horizons $\mathrm{H}$ in some black hole spacetimes of GR. For GR Kerr and GR Kerr-AdS (with its attractive cosmological constant), we have the progression $\mathcal{E}_{0}\rightarrow\mathrm{H}_{\mathrm{I}}\rightarrow\mathrm{H}_{\mathrm{O}}\rightarrow\mathcal{E}_{\mathrm{O}}$ at radii $r^{\mathcal{E}}_{0}<r^{\mathrm{H}}_{\mathrm{I}}<r^{\mathrm{H}}_{\mathrm{O}}<r^{\mathcal{E}}_{\mathrm{O}}$. Here, $\mathrm{H}_{\mathrm{I}}$ is the black hole’s inner horizon, while $\mathrm{H}_{\mathrm{O}}$ is its outer horizon. $\mathcal{E}_{\mathrm{O}}$ is thus the outer ergosurface surrounding the outer horizon.
@@
It is worthwhile to here discuss the transitions between dS and AdS spacetimes. We may think of GR Kerr black hole spacetimes as possessing a cosmological horizon $\mathrm{H}_{\mathrm{C}}$ and ergosurface $\mathcal{E}_{\mathrm{C}}$ at $r=+\infty$, making them asymptotically flat. When we have the repulsive cosmological constant $\Lambda>0$ of GR Kerr-dS spacetimes, these cosmological features are brought to finite $r$. For GR Kerr-AdS spacetimes, with $\Lambda<0$, these cosmological features are pushed out beyond $r=+\infty$.
@@ 3 Establishing causal structure
Correspondingly, in spacelike $\mathrm{S}$ spacetime regions $r$ is a timelike coordinate, so $g_{rr}<0$ and thus $\Delta^{\mathrm{H}}<0$. Within spacelike regions, motion may only go in one direction in $r$. These come in two varieties. In regions of the first type $\mathrm{S}^{-}$, only motion to decreasing $r$ is possible. This is what is commonly used to define black holes. In the second type $\mathrm{S}^{+}$, only motion to increasing $r$ is possible. This is now the type of spacelike region found past cosmological horizons.
To distinguish between an $\mathrm{S}^{-}$ and an $\mathrm{S}^{+}$ region, once we know that $\Delta^{\mathrm{H}}<0$, we look at the shape of the radial effective potential, and see whether trajectories are being attracted to smaller $r$ $(\mathrm{S}^{-})$ or repelled to larger $r$ $(\mathrm{S}^{+})$.
3.1 Radial Effective Potential
@@
where the dot represents differentiation with respect to to an affine parameter $\sigma$, and $p_{\theta}$ is the $\theta$-component of the test particle four-momentum $p_{\mu}$. The other new quantities $(\delta_{1},E,L_{z},\mathcal{Q})$ derive from four corresponding conservation relations. The constancy of the norm of a particle’s four-velocity $u^{\mu}$ may be expressed as
@@
Exploiting axisymmetry, we confine our derivation of a radial effective potential to the equatorial plane $(\theta=\frac{\uppi}{2})$, such that $p_{\theta}=0,\rho^{2}=r^{2},\Delta^{\theta}=1,$ and $\widetilde{Q}=0$. The radial equation of motion may then be written out as
@@
Following the discussion of [53], we see that this equation presents us with a reality condition encoding the fact that test particle motion is only possible where $\widetilde{R}(r)>0$. From this condition, we may solve for the energy $E$ of a massive test particle, where $\delta_{1}=1$. We get two solutions $E\geq V_{+}(r)$ and $E\leq V_{-}(r)$, where
@@
(b) $-V_{+}^{\prime}(r)$
Figure 1: Plots of $V_{+}(r)$ and $-V_{+}^{\prime}(r)$, for $(\gamma,\kappa,a,L_{z})=(0.1,0.01,0.25,0)$, within a timelike $\mathrm{T}$ region. We have an $\mathrm{S}^{-}$ region to the left of the event horizon at $r^{\mathrm{H}}_{\mathrm{O}}$, and an $\mathrm{S}^{+}$ region to the right of the cosmological horizon at $r^{\mathrm{H}}_{\mathrm{C}}$. The points (red) mark the values of these functions at the turning point of $V(r)$ at $r=r^{\mathrm{TP}}$.
[…]
Due to the presence of the square root of $\Upsilon(r;a,L_{z})$ in (14), $-V^{\prime}_{+}(r)$ is not well-defined within $\mathrm{S}$ regions, where $\Delta^{\mathrm{H}}<0$, and $-V^{\prime}_{+}(r)$ becomes complex. This then ensures that any turning points of $V_{+}(r)$, where $V^{\prime}_{+}(r^{\mathrm{TP}})=0$, must occur within $\mathrm{T}$ regions. Figure 1 shows representative plots of $V_{+}(r)$ and $-V_{+}^{\prime}(r)$ for parameter values $(\gamma,\kappa,a,L_{z})=(0.1,0.01,0.25,0)$.
If the horizon bounding an $\mathrm{S}$ region lies to the right of a turning point $(r^{\mathrm{TP}}<r^{\mathrm{H}})$, we calculate the sign of $-V_{+}^{\prime}(r)$ at some $r$ where $r^{\mathrm{TP}}<r<r^{\mathrm{H}}$. Likewise, if the horizon lies to the left of a turning point $(r^{\mathrm{H}}<r^{\mathrm{TP}})$, then we calculate the sign of $-V_{+}^{\prime}(r)$ at some $r$ where $r^{\mathrm{H}}<r<r^{\mathrm{TP}}$. If $-V_{+}^{\prime}(r)>0$ ($-V_{+}^{\prime}(r)<0$), then a test particle is repelled (attracted), and we have an $\mathrm{S}^{+}$ ($\mathrm{S}^{-}$) region.
This process of classifying $\mathrm{S}$ regions is clear from figure 1(b). Here, we see that $-V_{+}^{\prime}(r)>0$ to the right of the turning point $(r^{\mathrm{TP}}<r<r^{\mathrm{H}}_{\mathrm{C}})$, making the region immediately to the right of $r^{\mathrm{H}}_{\mathrm{C}}$ be $\mathrm{S}^{+}$. Correspondingly, $-V_{+}^{\prime}(r)<0$ to the left of the turning point $(r^{\mathrm{H}}_{\mathrm{O}}<r<r^{\mathrm{TP}})$, so the region immediately to the left of $r^{\mathrm{H}}_{\mathrm{O}}$ is $\mathrm{S}^{-}$.

### v2

@@ Abstract
We analyze the stationary, uncharged, rotating, vacuum solution to Weyl conformal gravity. We elucidate the causal and ergoregion structure of the spacetimes found in the parameter space of the metric for positive mass. These are then compared to the analogous Kerr, Kerr-de Sitter, and Kerr-Anti-de Sitter solutions to general relativity. Additionally, we investigate and derive the extremal limits for both the horizons and ergosurfaces of the spacetimes. The horizon surface gravities and Hawking temperatures at the extremal horizon limits are then calculated to show that they vanish.
@@ 1 Introduction
where $G_{\mu\nu}$ is the Einstein curvature tensor, and $T_{\mu\nu}$ is the stress-energy tensor. We use geometrized units, where $G=c=\hbar=1$, and a metric $g_{\mu\nu}$ signature of (,+,+,+)(-,+,+,+).
@@
[…]
The desire to resolve these tensions at astrophysical and cosmological scales motivates work into modified gravity theories such as Modified Newtonian dynamics (MOND) [17], massive gravity [18], Horndeski gravity [19, 20], and others.
It is also understood that GR breaks down at quantum mechanical scales. Attempting to use GR as a quantum field theory leads to non-renormalizable divergences at high energies [21]. The failure of GR in the miscroscopic regime is evident in the presence of spacetime singularities, with divergent curvature scalars, in black hole models and at the dawn of the universe [22].
The fundamental failings of GR at these ultraviolet (UV) scales calls for the formulation of viable quantum gravity theories. Current lines of work in this field include string theory [23], loop quantum gravity [24], twistor theory [25], and a plethora of others.
@@
Requiring conformal invariance in the field equations points to a gravitational action of the form
$I_{\rm W}=-\alpha_{g}\int\text{d}^{4}x\sqrt{-g}\ C_{\lambda\mu\nu\kappa}C^{\lambda\mu\nu\kappa},$ (2)
where $\alpha_{\text{g}}$ is a dimensionless gravitational coupling constant, $g$ is the determinant of the metric $g_{\mu\nu}$, and $C_{\lambda\mu\nu\kappa}$ is the Weyl tensor given by the traceless part of the Riemann tensor $R_{\lambda\mu\nu\kappa}$ [29]. This can be obtained by subtracting all traces from the Riemann tensor
$C_{\lambda\mu\nu\kappa}=R_{\lambda\mu\nu\kappa}$- $\dfrac{1}{2}(g_{\lambda\nu}R_{\mu\kappa}-g_{\lambda\kappa}R_{\mu\nu}-g_{\mu\nu}R_{\lambda\kappa}+g_{\mu\kappa}R_{\lambda\nu})$
$+\dfrac{1}{6}{R^{\alpha}}_{\alpha}(g_{\lambda\nu}g_{\mu\kappa}-g_{\lambda\kappa}g_{\mu\nu}),$ (3)
where, of course, $R_{\mu\nu}=g^{\rho\sigma}R_{\sigma\mu\rho\nu}$ is the Ricci tensor.
Varying the gravitational action $I_{\rm W}$ (2) with respect to the metric $g_{\mu\nu}$ gives [30]
$\frac{1}{\sqrt{-g}}\frac{\delta I_{\rm W}}{\delta g_{\mu\nu}}=-2\alpha_{g}W^{\mu\nu},$ (4)
where $W^{\mu\nu}$ is the traceless Bach tensor [31]. This may be written as
$W_{\mu\nu}=W^{(2)}_{\mu\nu}-\dfrac{1}{3}W^{(1)}_{\mu\nu}$, (5)
with
[…]
as given in [29].
This then allows us to write the CG field equations in a similar form to those for GR in (1) as
$W_{\mu\nu}=\frac{1}{4\ \alpha_{\text{g}}}\ T_{\mu\nu}.$ (7)
While $G_{\mu\nu}$ contains up to second-order derivatives of the metric, $W_{\mu\nu}$ has up to fourth-order derivatives, making CG a fourth-order theory [28, 31].
@@
$\rmd s^{2}=-B(r)\ \rmd t^{2}+\frac{\rmd r^{2}}{B(r)}+r^{2}\left(\mathrm{\leavevmode\nobreak\ d}\theta^{2}+\sin^{2}\theta\ \rmd\phi^{2}\right).$ (8)
@@
$B(r)=1-3\beta\gamma-\frac{2\widetilde{M}}{r}+\gamma r-\kappa r^{2}.$ (10)
@@
[…]
Conformal gravity successfully replicates results of GR in such classical tests as perihelion precession and time dilation [40, 41] for small values of $\gamma$ and $\kappa$, making it a viable alternative to GR.
Turning from these astrophysical and cosmological tensions to the UV regime, CG has also been considered as a candidate quantum gravity theory. Within the theory, for instance, conformal transformations have been shown to remove the divergence of curvature scalars at black hole singularities [27, 42]. Furthermore, the fact that the coupling constant $\alpha_{g}$ in the action $I_{\mathrm{W}}$ (2) is dimensionless allows CG to be power counting renormalizable [15, 30], without necessitating higher dimensions like string theory requires.
One of the main criticisms of CG, as with other fourth-order theories, is that it should be subject to ghosts. However, it has been shown that a proper formulation of CG as a quantum theory is, in fact, ghostless [43, 44].
Despite these successes, there remain open problems in CG. These range from the getting the correct proportion of deuterium in big bang nucleosynthesis [45], properly modelling galaxy cluster dynamics [46, 47], to computing the decay process of binary pulsar orbits [15].
@@
$\rmd s^{2}$ $=-\left(1-\frac{2\widetilde{M}r}{\rho^{2}}-k\left(r^{2}-a^{2}\cos^{2}\theta\right)\right)\ \rmd t^{2}$ (11)
$+2\left(\frac{-2\widetilde{M}ra\sin^{2}\theta+ka\left(a^{2}\left(r^{2}+a^{2}\right)\cos^{4}\theta-r^{4}\sin^{2}\theta\right)}{\rho^{2}}\right)\ \rmd t\ \rmd\phi$
$+\left(\frac{\rho^{2}}{\Delta^{\mathrm{H}}}\right)\rmd r^{2}+\left(\frac{\rho^{2}}{\Delta^{\theta}}\right)\ \rmd\theta^{2}+\left(\frac{\Sigma^{2}\sin^{2}\theta}{\rho^{2}}\right)\ \rmd\phi^{2}\ ,$
@@
From this, we recover the GR Kerr metric [53] for $\gamma,\kappa\rightarrow 0$. An appropriate conformal transformation and taking $\gamma\rightarrow 0$ reduces CG Kerr to GR Kerr-de Sitter (Kerr-dS) for $\kappa>0$, and GR Kerr-Anti-de Sitter (Kerr-AdS) for $\kappa<0$. Conformal transformations also show that CG Kerr (11) is equivalent to CG Schwarzschild (8) when $a\rightarrow 0$ [32, 51]. As such a transformation has not been explicitly shown in Boyer-Linquist coordinates in existing literature, we demonstrate this in A.
@@
We also calculate the surface gravities and the temperatures associated with Hawking radiation [56, 57] for the horizons in these spacetimes. We check and confirm that these quantities indeed vanish at the extremal horizon limits, as one would expect from the GR counterparts to the CG Kerr solution. Explorations of the horizon thermodynamics of some other CG metrics have been done by [58, 59].
In this work, we explore the parameter space of the CG parameters $\gamma$ and $\kappa$, the mass $\beta$, and the spin parameter $a$. We consider both positive and negative values of $\gamma$ and $\kappa$.
Values of $\gamma<0$ had previously been said to be required to permit gravitational lensing [41, 60], in contrast to $\gamma>0$ values used to fit galaxy rotation curves and cosmological phenomena [28, 30, 38]. However, more recent work [61, 62] has pointed out that erroneous approximations in calculating the bending angle in these previous studies of the lensing problem may have led to the conclusion that the wrong $(\gamma<0)$ sign was needed. Thus, $\gamma>0$ values have been found to align with both lensing and galactic rotation curve fits.
As we mentioned earlier, research to establish values for $\gamma$ and $\kappa$ by performing fits to rotation curves and cosmological expansion is still ongoing. Summarized in [32, 39], work by [30, 28, 38] give the orders of magnitude $\gamma\sim 10^{-30}-10^{-28}\ \text{cm}^{-1}$ and $\kappa\sim 10^{-54}-10^{-48}\ \text{cm}^{-2}$ when considering stars within galaxies as their sources. The solar system scale tests of perihelion precession [60], time delay, and light deflection [40, 41] fall within the magnitudes of these constraints. However, the specific values of such parameters may not necessarily be universal, and instead depend upon the details of the density distributions and structures of the sources being considered [33]. Thus, as we are considering theoretical black holes in vacuum metrics in this study, we do not confine ourselves to any particular range of $\gamma$ and $\kappa$ values.
Since we are most interested in black hole solutions, we confine ourselves to positive mass $\beta>0$. As a stationary metric, we have an invariance to a simultaneous transformation $\rmd t\rightarrow-\rmd t$ and $\rmd\phi\rightarrow-\rmd\phi$, so may then simply consider spins $a>0$.
@@
We begin, in section 2, by introducing the equations governing the presence of ergosurfaces and horizons. Afterwards, in section 3, we establish the means to determine the causal structure and nature of horizons and ergosurfaces. The relevant equations for the extremal limits of CG Kerr spacetimes are derived in section 4, and we also briefly explore the surface gravities and Hawking temperatures at these limits. We then move on, in section 5, to establishing the general structure of CG Kerr spacetimes for $\kappa=0$ along with elucidating the role of $\gamma$. The effects of varying $\kappa$ and the spin $a$ are handled in sections 6 and 7 respectively. We summarize our conclusions in section 8.
@@ 2.1 Ergoregions and ergosurfaces
The ergosurfaces, where $g_{tt}=0$, that bound ergoregions are then known as static limits. On the equatorial plane, where $\theta=\frac{\uppi}{2}$, this condition is given by
$\Delta^{\mathcal{E}}\equiv\left(-\rho^{2}g_{tt}\right)\big{|}_{\theta=\frac{\uppi}{2}}=-kr^{4}+r^{2}-2\widetilde{M}r=0.$ (13)
Since $\cos\frac{\uppi}{2}=0$, the spin $a$ dependent term in $g_{tt}$ (11) vanishes leaving $\Delta^{\mathcal{E}}$ with no explicit dependence on the spin $a$.
@@
As a quartic with real coefficients, (13) has four, two, or no real roots. Similar to GR, the CG Kerr solution possesses a ring singularity at $(r,\theta)=\left(0,\frac{\uppi}{2}\right)$. While real roots of $\Delta^{\mathcal{E}}$ may be found for $r<0$, we confine our work to discussions of features found for $r\geq 0$. A maximum of three ergosurfaces are found for $r\geq 0$.
@@ 2.2 Horizons
In the Boyer-Lindquist coordinates that we use in (11), horizons or static limits, meanwhile, are defined by where normal vectors $n_{\mu}$ to $r=\text{constant}$ surfaces become null. These occur when [67]
$n_{\mu}n_{\nu}\ g^{\mu\nu}=g^{rr}=\frac{\Delta^{\mathrm{H}}}{\rho^{2}}=0,$ (14)
where $\Delta^{\mathrm{H}}$ is as defined in (12). As we can see, this is equivalent to finding where $g_{rr}$ switches sign as $1/g_{rr}=\Delta^{\mathrm{H}}/\rho^{2}=0$.
This switching of sign of $g_{rr}$ means that $r$ transitions from being a spacelike coordinate $(g_{rr}>0)$ to a timelike one $(g_{rr}<0)$. In finding such horizons then, we seek to solve
@@
To better orient ourselves, we review the ordering of ergosurfaces $\mathcal{E}$ and horizons $\mathrm{H}$ in some black hole spacetimes of GR. For GR Kerr and GR Kerr-AdS (with its attractive cosmological constant), we have the progression $\mathcal{E}_{0}\rightarrow\mathrm{H}_{\mathrm{I}}\rightarrow\mathrm{H}_{\mathrm{O}}\rightarrow\mathcal{E}_{\mathrm{O}}$ at radii $r^{\mathcal{E}}_{0}<r^{\mathrm{H}}_{\mathrm{I}}<r^{\mathrm{H}}_{\mathrm{O}}<r^{\mathcal{E}}_{\mathrm{O}}$. Here, $\mathrm{H}_{\mathrm{I}}$ is the black hole’s inner horizon, while $\mathrm{H}_{\mathrm{O}}$ is its outer horizon. $\mathcal{E}_{\mathrm{O}}$ is thus the outer ergosurface surrounding the outer horizon.
@@
It is worthwhile to here discuss the transitions between dS and AdS spacetimes. We may think of GR Kerr black hole spacetimes as possessing a cosmological horizon $\mathrm{H}_{\mathrm{C}}$ and ergosurface $\mathcal{E}_{\mathrm{C}}$ at $r=+\infty$, making them asymptotically flat. When we have the repulsive cosmological constant $\Lambda>0$ of GR Kerr-dS spacetimes, these cosmological features ($\mathrm{H}_{\mathrm{C}}$ and $\mathcal{E}_{\mathrm{C}}$) are brought to finite $r$. For GR Kerr-AdS spacetimes, with $\Lambda<0$, these cosmological features are pushed out beyond $r=+\infty$.
@@ 3 Establishing causal structure
Correspondingly, in spacelike $\mathrm{S}$ regions $r$ is a timelike coordinate, so $g_{rr}<0$ and thus $\Delta^{\mathrm{H}}<0$. Within spacelike regions, world lines progress in only one direction in $r$. These come in two varieties. In regions of the first type $\mathrm{S}^{-}$, all world lines progress toward decreasing $r$, as in the interiors of GR Schwarzschild black holes. In the second type $\mathrm{S}^{+}$, motion is possible only toward increasing $r$. This is now the type of spacelike region found exterior (at greater $r$) to cosmological horizons.
To distinguish between an $\mathrm{S}^{-}$ and an $\mathrm{S}^{+}$ region, once we know that $\Delta^{\mathrm{H}}<0$, we look at the shape of the effective radial potential, and see whether trajectories are attracted to smaller $r$ $(\mathrm{S}^{-})$ or repelled to larger $r$ $(\mathrm{S}^{+})$.
3.1 Effective radial potential
@@
where the dot represents differentiation with respect to an affine parameter $\sigma$, and $p_{\theta}$ is the $\theta$-component of the test particle four-momentum $p_{\mu}$. The other new quantities $(\delta_{1},E,L_{z},\mathcal{Q})$ derive from four corresponding conservation relations. The constancy of the norm of a particle’s four-velocity $u^{\mu}$ may be expressed as
@@
Exploiting axisymmetry, we confine our derivation of a radial effective potential to the equatorial plane $(\theta=\frac{\uppi}{2})$, such that $p_{\theta}=0,\rho^{2}=r^{2},\Delta^{\theta}=1,$ and $\mathcal{Q}=0$. The radial equation of motion may then be written out as
@@
Following the discussion of [69], we see that this equation presents us with a reality condition encoding the fact that test particle motion is only possible where $\widetilde{R}(r)\geq 0$. From this condition, we may solve for the energy $E$ of a massive test particle, where $\delta_{1}=1$. We get two solutions $E\geq V_{+}(r)$ and $E\leq V_{-}(r)$, where
@@
(b) $V_{+}^{\prime}(r)$
Figure 1: Plots of $V_{+}(r)$ and $V_{+}^{\prime}(r)$, for $(\gamma,\kappa,a,L_{z})=(0.1,0.01,0.25,0)$, within a timelike $\mathrm{T}$ region. We have an $\mathrm{S}^{-}$ region to the left of the event horizon at $r^{\mathrm{H}}_{\mathrm{O}}$, and an $\mathrm{S}^{+}$ region to the right of the cosmological horizon at $r^{\mathrm{H}}_{\mathrm{C}}$. The points (red) mark the values of these functions at the turning point of $V(r)$ at $r=r^{\mathrm{TP}}$.
As we are considering massive test particles, we may take the affine parameter to be the proper time, such that $\sigma\rightarrow\tau$. For future-directed trajectories $(\rmd t/\rmd\tau>0)$, the positive root $V_{+}(r)$ is the relevant one to be used [69, 70].
To see whether a particle is attracted or repelled, we calculate the sign of $\ddot{r}=-V_{+}^{\prime}(r)\equiv-\rmd V_{+}(r)/\rmd r$. As our purpose here is merely to distinguish between $\mathrm{S}^{-}$ and $\mathrm{S}^{+}$ regions, for additional ease, we set a vanishing angular momentum $(L_{z}=0)$. As it is rather unwieldy, we do not write out $V^{\prime}(r)$ in full here.
Due to the presence of the square root of $\Upsilon(r;a,L_{z})$ in (20), $V^{\prime}_{+}(r)$ is not well-defined within $\mathrm{S}$ regions, where $\Delta^{\mathrm{H}}<0$, and $V^{\prime}_{+}(r)$ becomes complex. This then ensures that any turning points of $V_{+}(r)$, where $V^{\prime}_{+}(r^{\mathrm{TP}})=0$, must occur within $\mathrm{T}$ regions. Figure 1 shows representative plots of $V_{+}(r)$ and $V_{+}^{\prime}(r)$ for parameter values $(\gamma,\kappa,a,L_{z})=(0.1,0.01,0.25,0)$.
If the horizon bounding an $\mathrm{S}$ region lies to the right of a turning point $(r^{\mathrm{TP}}<r^{\mathrm{H}})$, we calculate the sign of $V_{+}^{\prime}(r)$ at some $r$ where $r^{\mathrm{TP}}<r<r^{\mathrm{H}}$. Likewise, if the horizon lies to the left of a turning point $(r^{\mathrm{H}}<r^{\mathrm{TP}})$, then we calculate the sign of $V_{+}^{\prime}(r)$ at some $r$ where $r^{\mathrm{H}}<r<r^{\mathrm{TP}}$. If $V_{+}^{\prime}(r)>0$ ($V_{+}^{\prime}(r)<0$), then a test particle is repelled (attracted), and we have an $\mathrm{S}^{+}$ ($\mathrm{S}^{-}$) region.
This process of classifying $\mathrm{S}$ regions is clear from figure 1(b). Here, we see that $V_{+}^{\prime}(r)>0$ to the right of the turning point $(r^{\mathrm{TP}}<r<r^{\mathrm{H}}_{\mathrm{C}})$, making the region immediately to the right of $r^{\mathrm{H}}_{\mathrm{C}}$ be $\mathrm{S}^{+}$. Correspondingly, $V_{+}^{\prime}(r)<0$ to the left of the turning point $(r^{\mathrm{H}}_{\mathrm{O}}<r<r^{\mathrm{TP}})$, so the region immediately to the left of $r^{\mathrm{H}}_{\mathrm{O}}$ is $\mathrm{S}^{-}$.